# AI red-teaming is a sociotechnical challenge: on values, labor, and harms

### v2

@@ Abstract.
As generative AI technologies find more and more real-world applications, the importance of testing their performance and safety seems paramount. “Red-teaming” has quickly become the primary approach to test AI models–prioritized by AI companies, and enshrined in AI policy and regulation. Members of red teams act as adversaries, probing AI systems to test their safety mechanisms and uncover vulnerabilities. Yet we know far too little about this work or its implications. This essay calls for collaboration between computer scientists and social scientists to study the sociotechnical systems surrounding AI technologies, including the work of red-teaming, to avoid repeating the mistakes of the recent past. We highlight the importance of understanding the values and assumptions behind red-teaming, the labor arrangements involved, and the psychological impacts on red-teamers, drawing insights from the lessons learned around the work of content moderation.
@@
Generative AI technology like Large Language Models (LLMs) and image diffusion models have emerged so rapidly, from research projects inside a few tech companies and university computer science departments, to the engines behind the global deployment of generative AI to consumers. Whether accessed directly on the web or embedded in software, search, or social media, generative AI is everywhere. When a technology jumps this quickly from experimental plaything to widely available consumer service, many other elements also settle in around it, often without much forethought: interfaces, policies, business models, labor arrangements, infrastructural assurances, complementary technologies, public claims, advertising campaigns, regulations.
Many of these decisions, arrangements, and infrastructures may turn out to be just as consequential for users and the broader public as the core technology itself. But the boisterous promises and debates that surround the new technology can obscure these other essential elements, that make technologies always more than the sum of their engineered parts. Researchers studying the workings and implications of these technologies, across computer science, engineering, the social sciences, humanities, and law, must gear up just as quickly, to study not just the core technology, but the sociotechnical system taking shape around it (Joyce et al., 2021). And they need to study it together.
@@
In this essay, we call upon computer scientists and social scientists to pay closer, critical, and collaborative attention to one part of AI development, “red-teaming.” [^1] The spelling of the term is inconsistent across different texts, organizations (and sometimes even within organizations). In line with (Friedler et al., 2023), we hyphenate the term except when specifically referencing the teams themselves. AI models and their applications typically undergo internal testing before release, and continue to be evaluated during use; red-teaming aims to probe these applications for exploitable vulnerabilities, hallucinations, and bias. From an AI evaluation vantage point, red-teaming is an assessment process that needs to be well-designed, effective, and replicable. But from a sociological vantage point, red-teaming is something else as well: a specific kind of labor, done by specific sets of people, in specific institutional contexts, with its own specific set of implications (Friedler et al., 2023; Metcalf and Singh, 2023). Our aim is to shed light on these sociotechnical elements, and to call for more cross-disciplinary attention to this critical component of AI development.
@@
Since the commercial launch of ChatGPT, red-teaming has been quickly normalized as a step in the production and deployment of generative AI models. AI developers champion it as proof of their public responsibility, while regulators count on it as a bulwark preventing AI from inflicting social harms. But the public knows precious little about how this work is conducted, upon what values and assumptions it is based, who is enlisted to do it, or the psychological costs they bear. This was the case with content moderation: too long hidden from public and critical scrutiny, labor and well-being concerns were too long overlooked; and the opacity of these value judgments soon became a political liability for social media platforms. And, given that what little the public does know about AI red-teaming has come from Silicon Valley’s own promotional materials, and given that the changing political climate particularly in the U.S. may discourage such performances of responsibility, the public may soon know even less.
@@
This essay is not based on a particular empirical research effort, nor is it based on any information internal to Microsoft. Rather, we share a set of observations drawn from our recent studies and collaborations concerned with the labor behind Responsible AI (Gray and Suri, 2019; Zhang et al., 2024b, a), the politics of generative AI as a new media technology (Gillespie, 2024), and participatory approaches to AI development and governance (Suresh et al., 2024; Deng et al., 2024). [^2] Also see Microsoft Research’s Project Resolve: https://www.microsoft.com/en-us/research/project/project-resolve/ And, being trained in science and technology studies, labor, psychology and design, we feel it is important to contextualize AI red-teaming with longer histories of scholarship that reckon with the social construction of new technologies.
@@ 1. What is red-teaming?
The term migrated to the field of cybersecurity. Red teams were tasked with infiltrating information systems to simulate worst case scenarios, like the theft of sensitive information or hacks on infrastructure, that might lead to financial and operational disaster (Zenko, 2015). Red teams became an important element of systems security, that might even pay for themselves.333see https://www.wired.com/story/microsoft-ai-red-team/
@@
For generative AI, red-teaming often concerns the security of the foundation model, but it increasingly also means purposefully provoking the model to produce undesired or incorrect responses. Because users might either intentionally or inadvertently prompt an AI model to generate hateful, pornographic, vile, or biased responses, red teams attempt to do so first, serving as a kind of adversary to the intended or presumed use, preemptively tempting the AI model to say things it shouldn’t and documenting how the safety architectures meant to prevent such responses can be shored up as a result.
@@
Still, what constitutes AI red-teaming remains fuzzy, given that its social and organizational practices and structures are still forming. Its relationship to evaluation, social engineering, bug bounties, threat assessments, and penetration testing is still ”being discovered” (Singh et al., 2025). Even the spelling of the term itself is unsettled: sometimes it is hyphenated or concatenated, other times not. Other terms from cybersecurity and hacker lingo bear a family resemblance: Google describes their AI Red Team as “ethical hackers”, a nod to venerable “white hat” hackers-for-hire who hunt for technical insecurities. [^4] https://blog.google/technology/safety-security/googles-ai-red-team-the-ethical-hackers-making-ai-safer/
@@
Some assert that red-teaming is essential to AI, and the surest way to safeguard equitable and responsible AI development. Others worry that red-teaming and the guardrails it produces are a kind of “security theater,” more performative than substantive (Feffer et al., 2024), meant to obscure the reckless deployment of harmful technologies to the public. Others suggest that efforts to police AI models will hinder the true potential of the technology, ultimately leading to “woke AI”. [^5] On X, Elon Musk claimed a “woke AI” may eventually kill people: https://x.com/elonmusk/status/1768746706043035827
@@
But in blunt economic terms, the first AI company to successfully “tame” their generative AI products through such safeguards could well capture the market. Business clients are asking for AI-powered customer service chatbots that do not hallucinate [^6] In 2024, Air Canada was found liable for its customer support chatbot’s incorrect claim about a customer’s ticket: https://www.cbc.ca/news/canada/british-columbia/air-canada-chatbot-lawsuit-1.7116416 and productivity tools that behave consistently. [^7] Complaints about AI tools getting “lazy” surface from time to time: https://www.theguardian.com/commentisfree/2024/jan/12/chatgpt-problems-lazy, as well as concerns about generative AI systems degrading over time: https://www.nytimes.com/interactive/2024/08/26/upshot/ai-synthetic-data.html There are enormous financial, reputational, and regulatory incentives to make generative AI tools safe and value-transparent, pushing these companies to rapidly institute red teams—arguably, faster than researchers concerned about the labor politics of AI can follow.
@@
Red-teaming also offers a kind of reassurance, easing fears held by the public, governments, and financial stakeholders about the safety and performance of generative AI systems. New AI products are often touted as having been tested by red-teamers before a wider release. [^8] For example, when OpenAI announced its Sora video generation model in early 2024, they prominently featured their red-teaming efforts on its launch page. https://web.archive.org/web/20240215192216/https://openai.com/sora#safety The US federal government at one point adopted the language of red-teaming as an important assurance of the safety of AI systems. In October 2023, US President Biden issued an Executive Order that reinforced the importance of red-teaming in a proposed system of federal oversight over AI: “Any foundation model that poses a serious risk to national security, national economic security, or national public health and safety… must share the results of all red-team safety tests.” [^9] White House Executive Order (14110) on the Safe, Secure and Trustworthy Development and Use of Artificial Intelligence, 2023 That order has since been rescinded, but red-teaming will certainly appear in subsequent efforts to regulate, or self-regulate, generative AI.
@@
As a new labor formation, developing under financial and political pressure, red-teaming echoes other contingent forms of digital labor: data labeling, content moderation, enrichment services of all kinds–all arrangements for the semi-automation of human judgment critical to data-driven technical systems. If we want AI that is not only safe and secure but also sustainable, we need to also study the labor arrangements emerging that are critical to it: the mix of internal teams testing out AI products, volunteers convened at hack-a-thon-like events, third party crowdwork vendors, and professional security firms. And, as we learned in the case of social media, we must attend to the psychological costs of the work asked of red teams tasked with making AI safe, secure, and useful to everyone.
@@
Rather than focus exclusively on implementing and improving it, for the model’s sake, we need to better understand red-teaming as a practice, and understand its place in the development of generative AI tools. A sociological perspective, which is fundamentally human-centric, can guide the responsible and effective use of red teams.
@@ 2. Value Judgments
Efforts to insert obligations, structures, and benchmarks have had to race to keep up with the rush to commercialize generative AI models (Lazar and Nelson, 2023). Initial red-teaming efforts were being implemented even as design teams were still wondering which harms to even probe for: AI red-teamers have had to develop homegrown taxonomies of harms, and the measurement and benchmarking systems for mitigating them (Hao et al., 2023; Weidinger et al., 2023). In public statements, the major AI companies often take as given that generative AI tools will unavoidably produce harmful content (Marres et al., 2024). However, it is much rarer for them to discuss how they determine what counts as harmful content, what they should and should not be looking for, and whether their teams are best suited to make those judgments. This prompts the question, “whose values are being utilized for alignment and evaluation?” (Feffer et al., 2024)
@@
The development of AI red-teaming echoes the early days of commercial content moderation at social media platforms. [^10] https://www.techpolicy.press/ais-content-moderation-moment-is-here/ The parallels are revealing. The categories of concern are strikingly similar: graphic violence, hate speech, harassment, discrimination, sexual content, terrorism, human trafficking, self-harm, child abuse, and misinformation  (Ganguli et al., 2022; noa, 2023). And when Silicon Valley found itself compelled to manage the gap between what can be generated online and what users should actually see, it too enlisted contingent human labor to serve as that filter.
@@
Social media platforms “discovered” the need for human moderation labor after being surprised by the kinds of content that could turn up through their services (Gillespie, 2018; Klonick, 2018; Gray and Suri, 2019). This awareness often came from user complaints, the technology press calling out the platforms’ shortcomings, and stumbling upon it themselves. AI red-teaming has similarly been fueled by user complaints and critical press coverage. Solving these harms has to be done internally and largely in proprietary ways (Carmi, 2019; Roberts, 2016), making industry-wide or public-wide discussions about harms and values difficult to develop. Leaving the public out of this process leaves the value judgments to the AI designers themselves. Or as OpenAI explained, echoing so many social media companies before them, “Our approach is to red-team iteratively, starting with an initial hypothesis of which areas may be the highest risk, testing these areas, and adjusting as we go.” (noa, 2023)
@@
Tackling harmful content internally, intuitively, and iteratively had profound implications for social media platforms over the past two decades; the same implications could befall the red-teaming of generative AI systems. Like AI red teams today, many social media platforms turned first to their own engineers and employees to evaluate for harmful content; to them, some categories of harms tend to seem more obvious, others tend to go unrecognized. Silicon Valley engineers generally do not reflect the range of identities and contexts of their global user base; social media platforms that began with their own employees often underestimated the harms faced by women, marginalized racial and ethnic groups, and those with stigmatized sexual and gender identities  (Gerrard and Thornham, 2020; Roberts, 2016; Ruckenstein and Turunen, 2019; Shahid and Vashistha, 2023).
@@
In the earliest days of social media platforms, the overriding approach to moderation was the “Feel bad? Take it down” rule (Klonick, 2018). Subjective judgments and gut feelings stood in, sometimes poorly, for the public’s understandings of harms or ethical principles. AI red-teaming strategies will have to be more complex and more inclusive to avoid the mistakes of social media’s past. However, with little opportunity for outside researchers to study commercial red teams, and little internal or external incentive to disclose much about their practices, we lack a clear empirical understanding of the harms that are being under-attended to, or fall outside the purview of a commercial organization.
@@ 3. Labor politics
The tendency to ignore the importance of human labor in AI systems, whether intentional or not, is common (Gray and Suri, 2019; Roberts, 2019). To this point, it is illustrative that ‘red-teaming’ is often referred to as a verb, eliding the human workforce that constitutes red teams. A sociotechnical examination of red-teaming should extend not only to the values and concerns behind the techniques red teams deploy, but to the people doing the work and the labor arrangements within which they operate.
@@
Red-teaming as a method is emerging in various forms: inside and outside companies, salaried and volunteer, with access to the inner workings of the AI system and without. These are bound to change: some forms will fall away, while others will settle in as “the way things are done.” But whatever particular labor politics do settle into place, there are important questions, ripe for scholarly analysis, about the institutional contexts, material conditions, and economic incentives of this AI-related work.
@@
Who does this work internally can vary. Big tech companies are eager to boast about their flagship red teams, [^11] Google even produced a short documentary on their premier red-teamers that has been viewed over 1 million times as of 2025: https://youtu.be/TusQWn2TQxQ?si=UnHO87JYDk1dblvK whose jobs are dedicated solely to ethical hacking, but we know that red-teaming also happens at smaller scales throughout the product development cycle (Singh et al., 2025). People doing the red-teaming may be part of larger Responsible AI efforts, Trust & Safety divisions, or legal/compliance apparatuses (Zhang et al., 2024a). Those who perform red-teaming for their own companies typically enjoy the job security of full employment. That said, they may not be in a position to refuse a red-teaming request. They are likely to have the necessary technical understanding of how models work, but it is not clear that this is sufficient to effectively identify and mitigate AI risks (Singh et al., 2025). And they may not be able to raise concerns publicly without breaching corporate norms or legally binding non-disclosure agreements (Christin et al., 2024).
@@
Employee red-teamers may have little training in any other relevant proficiencies, whether linguistic, sociocultural, historical, legal, or ethical; the incentive structures do not ask for or reward this expertise. Some internal red teams might include someone with sociocultural domain expertise, but they, too, work for the company and may have conflicting incentives (Friedler et al., 2023). Those windows of opportunity can open, a little. To test GPT-4 ahead of its March 2023 release, OpenAI solicited help from more than 50 experts, though primarily from Trust and Safety and cybersecurity backgrounds (noa, 2023). Anthropic and Microsoft encourage their red teams to consult with experts to test specific types of harms (Microsoft, 2023). These partnerships give AI companies access to experts without having to retain them as formal employees, compensating them in clout, API credits, job opportunities, or bragging rights rather than dollars.
@@
Following a pattern ubiquitous in Silicon Valley, there are increasing efforts to shift red-teaming labor from company employees to third-party datawork vendors, often overseas. Early in their model development, researchers at Anthropic enlisted several hundred untrained crowdworkers and instructed them to “make the AI behave badly” to elicit harmful responses from their LLM chatbot (Ganguli et al., 2022), In this project, the crowdworkers were responsible for both probing the AI and assessing its responses for harmful content.
@@
And, of course, red teams are expensive. Smaller companies competing to bring their generative AI services to market may not be equipped to employ internal red-teaming services sufficient to satisfy internal liability concerns or the specter of regulatory obligation. They may turn to an emerging crop of boutique AI safety and data services [^12] Such as Noma, Hidden Layer, Protect AI and Mindgard that see market opportunities in sourcing red-teaming. As scholars have noted (Gray and Suri, 2019), when labor is offloaded, particularly piecemeal to a globally-distributed contingent workforce, it becomes harder to trace and trickier to assert labor protections for it.
@@
The event demonstrated the viability of convening a greater diversity of perspectives than is represented on most internal company red teams. AI ethicist Rumman Chowdhury has argued that including diverse perspectives and backgrounds is imperative when red-teaming for topics such as race, gender, sexuality, politics, and class (Oremus, 2023). However, asking marginalized communities to elaborate on their own marginalization for red-teaming can easily slip into a transactional, extractive, and exploitative codependency (Dalal et al., 2024). Volunteer and uncompensated work requires significant effort to organize and has no obvious scaling mechanism. Access to generative AI models must be negotiated with AI companies, and actionable insights depend on the companies’ goodwill.
In many of the major generative AI applications, end users can also provide feedback, at least in limited forms. ChatGPT’s “thumbs down” includes a pull-down menu for users to indicate their concern: “Don’t like the style,” “Not factually correct,” “Refused when it shouldn’t have,” and others. It is not evident how the feedback gets back to designers, putting users in the position of performing opaque slivers of digital “civic labor” (Matias, 2019). User feedback is not strictly speaking red-teaming, as it isn’t necessarily adversarial. But end users do, whether inadvertently, playfully, or deviously, help designers surface outputs that companies did not anticipate or design for.
Red-teaming is still taking shape as a set of labor practices, inside and outside of AI companies. It is unclear whether or not companies’ internal red-teaming efforts are limited by the diversity of their employee pool and whether they may lean unfairly on their own minoritized employees to inject some diversity into their models (Benjamin, 2019; Lazar and Nelson, 2023). And outsourced red-teaming work raises tragically familiar concerns, echoing social media companies’ use of human labor. Outsourcing data work and using labor arbitrage to reduce costs leaves workers with fewer labor protections, more adverse work conditions, and more precarious job security, with few or no avenues for career development.

### v3

# AI red-teaming is a sociotechnical problem: on values, labor, and harms

@@ Abstract.
As generative AI technologies find more and more real-world applications, the importance of testing their performance and safety is paramount. “Red-teaming” has quickly become the primary approach to testing AI models–prioritized by AI companies, and enshrined in AI policy and regulation. Members of red teams act as adversaries, probing AI systems to test their safety mechanisms and uncover vulnerabilities. Yet we know far too little about this work or its implications. In this essay we highlight the importance of understanding the values and assumptions behind red-teaming, the labor arrangements involved, and the psychological impacts on red-teamers, drawing insights from lessons learned around the work of content moderation. Red-teaming should be a deeply interdisciplinary concern. To avoid repeating the mistakes of the recent past, we call for a coordinated network of scholars, from the full range of the computational and social sciences, to study the technical, social, critical, and policy dimensions of red-teaming and of the emerging sociotechnical system that is AI.
@@
Large language models (LLMs) and image diffusion models have emerged so rapidly, from research projects into the engines behind the global deployment of generative AI, whether accessed directly on the web or embedded in software, search, or social media. When a technology jumps this quickly from experimental plaything to widely available consumer service, many other elements also settle in around it, often without much forethought: interfaces, policies, business models, labor arrangements, infrastructural assurances, complementary technologies, public claims, advertising campaigns, regulations. Many of these decisions, arrangements, and infrastructures may turn out to be just as consequential for users and the broader public as the core technology itself. But the boisterous promises and debates that surround the new technology can obscure these other essential elements, elements that make technologies always more than the sum of their engineered parts.
Researchers studying the workings and implications of these technologies, across computer science, engineering, the social sciences, humanities, and law, must gear up just as quickly, to study not just the core technology, but the sociotechnical system taking shape around it (Joyce et al., 2021). And they need to study it together. In this essay, we call upon computer scientists and social scientists alike to pay closer, critical, and collaborative attention to one part of AI development, “red-teaming.” [^1] The spelling of the term is inconsistent across different texts and organizations (sometimes even within organizations): sometimes it is hyphenated or concatenated, other times not. In line with (Friedler et al., 2023), we hyphenate the term except when specifically referencing the teams themselves.
@@
AI models and their applications typically undergo internal testing before release, and continue to be evaluated during use; one part of this testing, red-teaming, probes these applications for exploitable vulnerabilities, errors, and bias. From an AI evaluation vantage point, red-teaming needs to be well-designed, effective, and replicable. But from a sociological vantage point, red-teaming is something else as well: a specific kind of labor, done by specific sets of people, in specific institutional contexts, with its own specific set of implications (Friedler et al., 2023; Metcalf and Singh, 2023). Our aim is to shed light on these sociotechnical elements, and to call for more cross-disciplinary attention to this critical component of AI development.
@@
Since the commercial launch of ChatGPT, red-teaming has been quickly normalized as a step in the production and deployment of generative AI models. AI developers champion it as proof of their public responsibility, while regulators count on it as a bulwark preventing AI from inflicting social harms. But the public knows precious little about how this work is conducted, upon what values and assumptions it is based, who is enlisted to do it, or the psychological costs they bear. This was also the case with content moderation: too long hidden from public and critical scrutiny, its labor and well-being concerns were too long overlooked; and the opacity of these value judgments soon became a political liability for social media platforms. And, given that what little the public does know about AI red-teaming comes largely from Silicon Valley’s own promotional materials, and given that the changing political climate particularly in the U.S. may discourage such outward performances of responsibility, the public may soon know even less.
@@
It is worth noting that this essay is not based on any information internal to Microsoft. Rather, we share our observations, drawn from our recent studies of the labor behind Responsible AI (Gray and Suri, 2019; Zhang et al., 2024b, a), the politics of generative AI as a new media technology (Gillespie, 2024), and participatory approaches to AI development and governance (Suresh et al., 2024; Deng et al., 2024). [^2] Also see Microsoft Research’s Project Resolve: https://www.microsoft.com/en-us/research/project/project-resolve/ And, being trained in science and technology studies, labor, psychology and design, we feel it is important to contextualize AI red-teaming within longer histories of scholarship that reckon with the social construction of new technologies.
@@ 1. What is red-teaming?
The term migrated to the field of cybersecurity. Red teams were tasked with infiltrating information systems to simulate worst case scenarios, like the theft of sensitive information or hacks on infrastructure, that might lead to financial or operational disaster (Zenko, 2015). Red teams became an important element of systems security, that might even pay for themselves by anticipating and thereby preventing breaches.333see https://www.wired.com/story/microsoft-ai-red-team/
@@
For generative AI, red-teaming sometimes concerns the security of the foundation model, but increasingly it also means purposefully provoking the model to produce undesired or incorrect responses. Because users might either intentionally or inadvertently prompt an AI model to generate hateful, pornographic, vile, or biased responses, red teams attempt to do so first, acting as a kind of adversary to the intended or presumed use. By preemptively tempting the AI model to say things it shouldn’t, they can document how to shore up the safety architectures meant to prevent such responses.
@@
Still, what constitutes AI red-teaming remains fuzzy, given that its tactics and organizational structures are still forming. Its familial resemblance to evaluation, social engineering, bug bounties, threat assessments, and penetration testing is still ”being discovered” (Singh et al., 2025). Other terms from cybersecurity and hacker lingo describe much the same project: Google describes their AI Red Team as “ethical hackers”, a nod to venerable “white hat” hackers-for-hire who hunt for technical insecurities. [^4] https://blog.google/technology/safety-security/googles-ai-red-team-the-ethical-hackers-making-ai-safer/
@@
Some assert that red-teaming is essential to AI, and the surest way to safeguard equitable and responsible AI development. Others worry that red-teaming and the guardrails it produces are a kind of “security theater,” more performative than substantive (Feffer et al., 2024), meant to obscure the reckless deployment of harmful technologies to the public. Others suggest that such efforts to police AI models will hinder the true potential of the technology, ultimately leading to “woke AI”. [^5] On X, Elon Musk claimed a “woke AI” may eventually kill people: https://x.com/elonmusk/status/1768746706043035827
@@
But in blunt economic terms, the first AI company to successfully “tame” their generative AI products through such safeguards could well capture the market. Business clients are asking for AI-powered customer service chatbots that do not hallucinate [^6] In 2024, Air Canada was found liable for its customer support chatbot’s incorrect claim about a customer’s ticket: https://www.cbc.ca/news/canada/british-columbia/air-canada-chatbot-lawsuit-1.7116416 and productivity tools that behave consistently. [^7] Complaints about AI tools getting “lazy” surface from time to time: https://www.theguardian.com/commentisfree/2024/jan/12/chatgpt-problems-lazy, as well as concerns about generative AI systems degrading over time: https://www.nytimes.com/interactive/2024/08/26/upshot/ai-synthetic-data.html There are enormous financial, reputational, and regulatory incentives to make generative AI tools safe and value-transparent, pushing these companies to rapidly institute red teams—arguably, faster than researchers concerned about the politics of AI can follow.
@@
Red-teaming also offers a kind of reassurance, easing fears held by the public, governments, and financial stakeholders about the safety and performance of generative AI systems. New AI products are often touted as having been tested by red-teamers before a wider release. [^8] For example, when OpenAI announced its Sora video generation model in early 2024, they prominently featured their red-teaming efforts on its launch page. https://web.archive.org/web/20240215192216/https://openai.com/sora#safety The US federal government (briefly) adopted the language of red-teaming as an important assurance of the safety of AI systems. In October 2023, US President Biden issued an Executive Order that reinforced the importance of red-teaming as part of a proposed system of federal oversight over AI: “Any foundation model that poses a serious risk to national security, national economic security, or national public health and safety… must share the results of all red-team safety tests.” [^9] White House Executive Order (14110) on the Safe, Secure and Trustworthy Development and Use of Artificial Intelligence, 2023 That order has since been rescinded, but red-teaming will likely appear in subsequent efforts to regulate, or self-regulate, generative AI.
@@
As a new labor formation, developing under financial and political pressure, red-teaming echoes other contingent forms of digital labor: data labeling, content moderation, enrichment services of all kinds–all arrangements for supporting the semi-automation of human judgment critical to data-driven technical systems. If we want AI that is not only safe and secure but also sustainable, we need to also study the labor arrangements emerging that are critical to it: the mix of internal teams testing out AI products, volunteers convened at hack-a-thon-like events, third party crowdwork vendors, and professional security firms. And, as we learned in the case of social media, we must attend to the psychological costs of the work asked of red teams charged with making AI safe, secure, and useful to everyone.
@@
Rather than focus exclusively on implementing and improving it, for the sake of the model, we need to better understand red-teaming as a practice, and understand its place in the development of generative AI tools, for the sake of the people involved. A sociological perspective, which is fundamentally human-centric, can better guide the responsible and effective use of red teams.
@@ 2. Value Judgments
Efforts to insert obligations, structures, and benchmarks have had to race to keep up with the rush to commercialize generative AI models (Lazar and Nelson, 2023). Initial red-teaming efforts were being implemented even as design teams were still wondering which harms to even probe for: AI red-teamers have had to develop homegrown taxonomies of harms, and the measurement and benchmarking systems for mitigating them (Hao et al., 2023; Weidinger et al., 2023). In public statements, the major AI companies often state as a given that generative AI tools will unavoidably produce harmful content (Marres et al., 2024). However, it is much rarer for them to discuss how they determine what counts as harmful content, what they should and should not be looking for, and whether their own teams are best suited to make those judgments. This prompts the question, “whose values are being utilized for alignment and evaluation?” (Feffer et al., 2024)
@@
The development of AI red-teaming echoes the early days of commercial content moderation at social media platforms. [^10] https://www.techpolicy.press/ais-content-moderation-moment-is-here/ The parallels are revealing. The categories of concern are strikingly similar: graphic violence, hate speech, harassment, discrimination, sexual content, terrorism, human trafficking, self-harm, child abuse, and misinformation  (Ganguli et al., 2022; OpenAI, 2023). And when Silicon Valley found itself compelled to manage the gap between what can be generated online and what users should actually see, it too enlisted contingent human labor to serve as that filter.
@@
Social media platforms “discovered” the need for human moderation labor after being surprised by the kinds of content that could turn up through their services (Gillespie, 2018; Klonick, 2018; Gray and Suri, 2019). This awareness often came from user complaints, the technology press calling out the platforms’ shortcomings, and platform companies stumbling upon it themselves. AI red-teaming has similarly been fueled by user complaints and critical press coverage. Solving these harms has to be done internally and largely in proprietary ways (Carmi, 2019; Roberts, 2016), making industry-wide or public-wide discussions about harms and values difficult to develop. But leaving the public out of this process leaves the value judgments to the AI designers themselves. Or as OpenAI explained, echoing so many social media companies before them, “Our approach is to red-team iteratively, starting with an initial hypothesis of which areas may be the highest risk, testing these areas, and adjusting as we go.” (OpenAI, 2023)
@@
Tackling harmful content internally, intuitively, and iteratively had profound implications for social media platforms over the past two decades; the same implications could befall the red-teaming of generative AI systems. Like AI red teams today, many social media platforms turned first to their own engineers and employees to evaluate for harmful content; to them, some categories of harms tended to seem more obvious, others tended to go unrecognized. Silicon Valley engineers generally do not reflect the range of identities and contexts of their global user base; social media platforms that began with their own employees often underestimated the harms faced by women, marginalized racial and ethnic groups, and those with stigmatized sexual and gender identities  (Gerrard and Thornham, 2020; Roberts, 2016; Ruckenstein and Turunen, 2019; Shahid and Vashistha, 2023).
@@
In the earliest days of social media platforms, the predominant approach to moderation was the “Feel bad? Take it down” rule (Klonick, 2018). Subjective judgments and gut feelings stood in, often poorly, for the public’s understandings of harms or ethical principles. AI red-teaming strategies will have to be more complex and more inclusive to avoid the mistakes of social media’s past. However, with little opportunity for outside researchers to study commercial red teams, and little internal or external incentive to disclose much about their practices, we lack a clear empirical understanding of what harms may be under-attended to, or fall outside the purview of a commercial organization.
@@ 3. Labor politics
The failure to appreciate the importance of human labor in AI systems, whether intentional or not, is common (Gray and Suri, 2019; Roberts, 2019). To this point, it is illustrative that ‘red-teaming’ is often referred to as a verb, eliding the human workforce that constitutes red teams. A sociotechnical examination of red-teaming should extend not only to the values and concerns behind the techniques red teams deploy, but to the people doing the work and the labor arrangements within which they operate.
@@
Red-teaming as a method is emerging in various forms: inside and outside companies, salaried and volunteer, with access to the inner workings of the AI system and without. These are bound to change: some forms will fall away, while others will settle in as “the way things are done.” But whatever particular labor politics do settle into place, there are important questions about the institutional contexts, material conditions, and economic incentives of this AI-related work, ripe for scholarly analysis.
@@
Who does this work internally can vary. Big tech companies are eager to boast about their flagship red teams, [^11] Google even produced a short documentary on their premier red-teamers that has been viewed over 1.5 million times as of December 2025: https://youtu.be/TusQWn2TQxQ?si=UnHO87JYDk1dblvK whose jobs are dedicated solely to ethical hacking, but we know that red-teaming also happens at smaller scales throughout the product development cycle (Singh et al., 2025). People doing the red-teaming may be part of larger Responsible AI efforts, Trust & Safety divisions, or legal/compliance apparatuses (Zhang et al., 2024a). While those who perform red-teaming for their own companies typically enjoy the job security of full employment, they may not be in a position to refuse a red-teaming request. While they are likely to have the necessary technical understanding of how models work, it is not clear that this is sufficient to effectively identify and mitigate AI risks (Singh et al., 2025). And they may not be able to raise concerns publicly without breaching corporate norms or legally binding non-disclosure agreements (Christin et al., 2024).
@@
Employee red-teamers may have little training in any other relevant proficiencies, whether linguistic, sociocultural, historical, legal, or ethical; the incentive structures do not ask for or reward such expertise. Some internal red teams might include someone with sociocultural domain expertise, but they, too, work for the company and may have conflicting incentives (Friedler et al., 2023). Those windows of opportunity can open, a little. To test GPT-4 ahead of its March 2023 release, OpenAI solicited help from more than 50 experts, though primarily from Trust and Safety and cybersecurity backgrounds (OpenAI, 2023). Anthropic and Microsoft encourage their red teams to consult with experts to test specific types of harms (Microsoft, 2023). These partnerships give AI companies access to experts without having to retain them as formal employees, compensating them in clout, API credits, job opportunities, or bragging rights rather than dollars.
@@
Following a pattern ubiquitous in Silicon Valley, there have been increasing efforts to shift red-teaming labor from company employees to third-party datawork vendors, often overseas. Doing so uses labor arbitrage to drive down wages and, as the history of outsourcing and offshoring demostrates, can erode worker protections and their capacity to raise concerns about the effectiveness of their work. Early in their model development, researchers at Anthropic enlisted several hundred untrained crowdworkers and instructed them to “make the AI behave badly” to elicit harmful responses from their LLM chatbot (Ganguli et al., 2022), In this project, the crowdworkers were responsible for both probing the AI and assessing its responses for harmful content.
@@
Either way, red teams can be an expensive undertaking. Smaller companies competing to bring their generative AI services to market may not be equipped to employ internal red-teaming services sufficient to satisfy internal liability concerns or regulatory obligations. They may turn to an emerging crop of boutique AI safety and data services [^12] Such as Noma, Hidden Layer, Protect AI and Mindgard that see market opportunities in sourcing red-teaming. As scholars have noted (Gray and Suri, 2019), when labor is offloaded, particularly piecemeal to a globally-distributed contingent workforce, it becomes harder to trace and trickier to assert labor protections for it.
@@
The event demonstrated the viability of convening a greater diversity of perspectives than is represented on most internal company red teams, something AI ethicist Rumman Chowdhury has argued is imperative when red-teaming on issues such as race, gender, sexuality, politics, and class (Oremus, 2023). However, asking marginalized communities to elaborate on their own marginalization for red-teaming can easily slip into a transactional, extractive, and exploitative codependency (Dalal et al., 2024). Volunteer and uncompensated work also requires significant effort to organize and has no obvious scaling mechanism. Access to generative AI models must be negotiated with AI companies, and taking any action on the insights gleaned depends entirely on the companies’ goodwill.
In many of the major generative AI applications, end users can also provide feedback, at least in limited forms. ChatGPT’s “thumbs down” includes a pull-down menu for users to indicate their concern: “Don’t like the style,” “Not factually correct,” “Refused when it shouldn’t have,” and others. But it is not evident how feedback from these slivers of digital “civic labor” (Matias, 2019) ever gets back to designers. User feedback is not strictly speaking red-teaming, as it isn’t necessarily adversarial. But end users do, whether inadvertently, playfully, or deviously, help designers surface outputs that companies did not anticipate or design for.
Red-teaming is still taking shape as a set of labor practices, inside and outside of AI companies. It is unclear whether or not companies’ internal red-teaming efforts are limited by the diversity of their employee pool, and how often they lean unfairly on their own minoritized employees to inject some diversity into their models (Benjamin, 2019; Lazar and Nelson, 2023). And, outsourced red-teaming work and using labor arbitrage to reduce costs raises tragically familiar concerns that echo social media companies’ use of human labor: leaving workers with fewer labor protections, more adverse work conditions, and more precarious job security, with few or no avenues for career development.